Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Three men and 2 women will present 5 consecutive speeches

Expert replies
by BTGmoderatorLU » Thu May 24, 2018 2:59 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Three men and 2 women will present 5 consecutive speeches, 1 by each person, at a conference. If the order of the speakers is determined randomly, what is the probability that at least 2 of the men's speeches will be consecutive?

A. 3124/3125
B. 9/10
C. 4/5
D. 16/25
E. 1/2

The OA is B.

Please, can anyone assist me with this PS question? How can I solve it? Thanks in advance!
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Thu May 24, 2018 3:54 pm
BTGmoderatorLU wrote:Three men and 2 women will present 5 consecutive speeches, 1 by each person, at a conference. If the order of the speakers is determined randomly, what is the probability that at least 2 of the men's speeches will be consecutive?

A. 3124/3125
B. 9/10
C. 4/5
D. 16/25
E. 1/2
P(at least 2 men speak consecutively) = 1 - P(no men speak consecutively).

For no men to speak consecutively, the order of the speeches must be as follows:
man-woman-man-woman-man.
P(1st speaker is a man) = 3/5. (Of the 5 speakers, 3 are men.)
P(2nd speaker is a woman) = 2/4. (Of the 4 remaining speakers, 2 are women.)
P(3rd speaker is a man) = 2/3. (Of the 3 remaining speakers, 2 are men.)
P(4th speaker is a woman) = 1/2. (Of the 2 remaining speakers, 1 is a woman.)
P(5th speaker is a man) = 1/1. (Only 1 speaker remains, and he is a man.)
To combine the probabilities above, we multiply:
P(no men speak consecutively) = 3/5 * 2/4 * 2/3 * 1/2 * 1/1 = 1/10.

Thus:
P(at least 2 men speak consecutively) = 1 - 1/10 = 9/10.

The correct answer is B.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by [email protected] » Sun May 27, 2018 6:41 pm
Hi All,

We're told that 3 men and 2 women will present 5 consecutive speeches (1 by each person) at a conference and that the order of the speakers is determined randomly. We're asked for the probability that at least 2 of the men's speeches will be consecutive. This question can be solved in a couple of different ways, including by 'brute force' (simply 'drawing out' the possibilities).

With 3 men and 2 women, the following orders are possible:

MMMWW
MMWMW
MWMMW
WMMMW

MMWWM
MWMWM
WMMWM

MWWMM
WMWMM

WWMMM

There are just 10 possibilities (and they are all equally likely). Of these 10 options, 9 of the 10 include at least 2 consecutive speeches by men.

Final Answer: B

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Tue May 29, 2018 6:38 am
BTGmoderatorLU wrote:Three men and 2 women will present 5 consecutive speeches, 1 by each person, at a conference. If the order of the speakers is determined randomly, what is the probability that at least 2 of the men's speeches will be consecutive?

A. 3124/3125
B. 9/10
C. 4/5
D. 16/25
E. 1/2
Since there are 5 people, there are 5! = 120 ways the speeches can be presented. The only way the 2 of the men’s speeches will not be consecutive is if the order is (M = a man’s speech and W = a woman’s speech): M-W-M-W-M. The number of ways this can be done is 3 x 2 x 2 x 1 x 1 = 12. So there must be 120 - 12 = 108 ways that at least two of the men’s speeches will be consecutive and thus the probability is 108/120 = 9/10.

Alternate Solution:

Using the formula for permutations with indistinguishable objects, 3 men and 2 women can present speeches in 5!/(3! x 2!) = (5 x 4)/(2 x 1) = 10 ways.

Of these 10 orderings of speeches, only M-W-M-W-M does not meet the requirement of having at least two consecutive talks by men; thus, 10 - 1 = 9 orderings do meet this requirement. Thus, the probability is 9/10.

Answer: B

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion